Question regarding the below example.
E and F are mutually exclusive and P(E) = 0.2 and P(F) = 0.5. Find P(E
U F).
If the two are mutually exclusive, using the addition rule, wouldn't P(E U F)=0.7? This is just a guess on my part as I don't know how to work through this type of problem.
Can someone please provide an explanation?
Thank you!

Found 2 solutions by ikleyn, MathLover1:
Answer by ikleyn(52781)

.
If the two are mutually exclusive, it means that P(E & F) = 0,


and then, using the general formula, you have


    P(E U F) = P(E) + P(F) - P(E & F) = P(E) + P(F) - 0 = P(E) + P(F).


So, your suggestion is correct.


The other, more visual way to see it is  to notice that if the sets E and F are mutually exclusive, then the sets E and F are disjoint.



Answer by MathLover1(20849)


E and F are mutually exclusive and P(E) = 0.2 and P(F) = 0.5. Find P(E
U F).
To calculate the probability that either E or F occurs we evaluate
P(E∪F) =P(E) +P(F) = 0.2+0.5 = 0.7